# Restricted Overpartitions and concave compositions: their modularity and asymptotics

### v1

@@ Abstract.
In this paper we study restricted overpartitions and concave compositions. Using $q$-series transformations, we show that their generating functions are related to modular forms, mock theta functions, false theta functions, and mock Maass theta functions. Moreover, we obtain their asymptotic main terms. We also study related rank statistics.
@@ 1. Introduction and Statements of results
Note that f(q)=R(1;q)f(q)=R(-1;q) is one of the mock theta functions introduced by Ramanujan in his last letter to Hardy. For Ramanujan [14], a mock theta function is a function $F$ that has asymptotic expansions at every roots of unity [^1] Equivalently, considering $F$ as a function of $\tau$, we can also consider asymptotic expansions at rationals. of the same type as those of weakly holomorphic modular forms, however there is not a single weakly holomorphic modular form whose asymptotic expansions agree at all roots of unity with that of $F$. In his thesis [38], Zwegers “completed” Ramanujan’s mock theta functions to harmonic Maass forms [^2] A harmonic Maass form is a function that satisfies modularity, is annihilated by the weight $k$ hyperbolic Laplacian, and satisfies a growth condition at the cusps. Its holomorphic part is called a mock modular form. by adding additional non-holomorphic terms. Ono and the second author [17] showed that, for roots of unity $\zeta\neq 1$, $R(\zeta;q)$ is the holomorphic part of a weight $\frac{1}{2}$ harmonic Maass form. Later Berkovich and Garvan [12, Section 5] defined the $M_{2}$-rank for partitions without repeated odd parts as the smallest integer at least $\frac{l(\lambda)}{2}$ minus the number of parts of $\lambda$, where $l(\lambda)$ is the largest part of $\lambda$. The $M_{2}$-rank generating function for partitions without repeated odd parts is given by
@@
False theta functions are theta functions but twisted with sign-factors. Due to this “twist”, they are not modular. Now, we give an example of a sequence whose generating function involves a false theta function. A unimodal sequence of $n$ is a finite sequence of positive integers $a_{j}~(1\leq j\leq r)$, $b_{\ell}~(1\leq\ell\leq s)$, $r,s\in\mathbb{N}_{0}$, and $\underline{c}\in\mathbb{N}$ such that $a_{r}\leq\ldots\leq a_{2}\leq a_{1}\leq\underline{c}\geq b_{1}\geq b_{2}\geq\ldots\geq b_{s}$ for some $k\in\mathbb{N}$ and $\sum_{j=1}^{r}a_{j}+\underline{c}+\sum_{j=1}^{s}b_{j}=n$. The generating function of [^3] In [3, Section 3], $u(n)$ was denoted by ${\sigma}{\sigma}(n)$ and was called stacks with summits of size $n$. $u(n)$, the number of unimodal sequences of size $n$, (see [3]) is given by
@@
Thus $U$ is a mixed false theta function, which a linear combinations of false theta functions multiplied by modular forms. Like Jacobi theta functions fit into the theory of Jacobi form, by introducing an extra Jacobi variable, Nazaroglu and the second author [20, Theorem 1.1] showed that the “modular completion” of a two-variable false theta functions transforms like a Jacobi form.
@@
In this paper, we consider overpartitions of $n$ with the following restrictions:
@@
Here we show that $\overline{P}_{\textup{od}}$ can be written in terms of a false theta function and a mixed mock Maass theta function [^5] Mixed mock Maass theta functions are linear combinations of products of modular forms and mock Maass theta functions.. For this define [^6] This is a mock Maass theta function by work of Zwegers (see the discussion at the end of Section 6 of [39]). Similar to the case of $\sigma$}, the Fourier coefficients of $W_{1}$ are related to the arithmetic of $\mathbb{Q}(\sqrt{2})$., following Corson, Favero, Leisinger, and Zubairy [24],
@@
Using this, [9, Theorem 6.1], Theorem 1.3, and the facts that $g(0)=0$ and $g(1)=1$, we obtain
@@
We recall another (see [21, p. 347]) generalized [^8] From [21, p. 347], we know that for $k,\ell\in\mathbb{N}$, $\nu(q^{k},q^{\ell};q)$ is a mock theta function in the sense of Ramanujan. third order mock theta function
@@ 2.1. $q$-series identities
Recall the following two identities from [5, equation (2.2.5) and equation (2.2.6)].
@@
Next, we require an identity due to Garvan [28, Theorem 1.2].
@@
Due to Kang [31, equation (7.5)], we have the following:
@@
In [25, Theorem 1.1], we find the following identity.
@@
From [6, equation (2.3.1)], we have the following lemma.
@@
From [5, Corollary 2.4], we record the following identity.
@@
From [5, Example 10, p. 29], we record the following identity.
@@
We also require identities by Bhoria, Eyyunni, and Maji. The first is (see [15, Theorem 2.1]).
@@
The second identity (see [15, Theorem 2.2]) is the following:
@@
Next, we list the following identity from [13, equation (3.2.6)].
@@
Finally, we require the Rogers–Fine identity [27, equation (14.1)].
@@ 2.2. A Tauberian theorem
We also require a modified version of Ingham’s Tauberian Theorem [30], stated in [19, Theorem 1.1]. For $\Delta\geq 0$ define
@@ 2.3. The Euler–Maclaurin summation formula
We use a variation of the Euler–Maclaurin summation formula, an important tool for computing asymptotic expansions of infinite sums, studied by Zagier [36, Proposition 3] for one-dimensional smooth functions and extended by Jennings-Shaffer, Mahlburg, and the second author [19, Theorem 1.2] for multi-dimensional complex-valued functions. A function $g$, defined over an unbounded domain $D\subset\mathbb{C}$, is of sufficient decay if there exists $\varepsilon>0$ such that $g(w)\ll|w|^{-1-\varepsilon}$ uniformly as $|w|\to\infty$ in $D$. For $0\leq\theta<\frac{\pi}{2}$, we define $D_{\theta}:=\{re^{i\alpha}:r\geq 0\ \text{and}\ |\alpha|\leq\theta\}$.
@@
Next, we consider two-dimensional sums. Throughout we write $\bm{x}=\left(x_{1},x_{2}\right)\in\mathbb{R}^{2}$. A function $f$ defined on an unbounded domain $D\subset\mathbb{C}^{2}$ is of sufficient decay if there exist $\varepsilon_{1},\varepsilon_{2}>0$ such that $f(\bm{x})\ll|w_{1}|^{-1-\varepsilon_{1}}|w_{2}|^{-1-\varepsilon_{2}}$ uniformly as $|w_{1}|+|w_{2}|\to\infty$ in $D$. From [19, Proposition 5.2] (with $r=2$), we have the following.
@@
From [24, Theorem 2.3], we have
@@
Plugging this into (5), using (1.9), using (1.4), [9, equation (1.6)], [27, equation (26.27)], and (1.13), we conclude the proof.∎
@@ 6. Concluding Remarks
We propose the following two questions for future research.

### v2

@@ Abstract.
In this paper we study restricted overpartitions and concave compositions. In several cases the resulting generating functions involve simultaneously modular forms, mock theta functions, mock Maass theta functions, and false theta functions, illustrating the appearance of mixed modular structures in restricted partition problems. Moreover, we obtain their asymptotic main terms. We also study related rank statistics.
@@ 1. Introduction and Statements of results
Note that f(q)=R(1;q)f(q)=R(-1;q) is one of the mock theta functions introduced by Ramanujan in his last letter to Hardy. For Ramanujan [15], a mock theta function is a function $F$ that has asymptotic expansions at every roots of unity [^1] Equivalently, considering $F$ as a function of $\tau$, we can also consider asymptotic expansions at rationals. of the same type as those of weakly holomorphic modular forms, however there is not a single weakly holomorphic modular form whose asymptotic expansions agree at all roots of unity with that of $F$. In his thesis [38], Zwegers “completed” Ramanujan’s mock theta functions to harmonic Maass forms [^2] A harmonic Maass form is a function that satisfies modularity, is annihilated by the weight $k$ hyperbolic Laplacian, and satisfies a growth condition at the cusps. Its holomorphic part is called a mock modular form. by adding additional non-holomorphic terms. Ono and the second author [18] showed that, for roots of unity $\zeta\neq 1$, $R(\zeta;q)$ is the holomorphic part of a weight $\frac{1}{2}$ harmonic Maass form. Later Berkovich and Garvan [13, Section 5] defined the $M_{2}$-rank for partitions without repeated odd parts as the smallest integer at least $\frac{l(\lambda)}{2}$ minus the number of parts of $\lambda$, where $l(\lambda)$ is the largest part of $\lambda$. The $M_{2}$-rank generating function for partitions without repeated odd parts is given by
@@
False theta functions are theta functions but twisted with sign-factors. Due to this “twist”, they are not modular. Now, we give an example of a sequence whose generating function involves a false theta function. A unimodal sequence of $n$ is a finite sequence of positive integers $a_{j}~(1\leq j\leq r)$, $b_{\ell}~(1\leq\ell\leq s)$, $r,s\in\mathbb{N}_{0}$, and $\underline{c}\in\mathbb{N}$ such that $a_{r}\leq\ldots\leq a_{2}\leq a_{1}\leq\underline{c}\geq b_{1}\geq b_{2}\geq\ldots\geq b_{s}$ for some $k\in\mathbb{N}$ and $\sum_{j=1}^{r}a_{j}+\underline{c}+\sum_{j=1}^{s}b_{j}=n$. The generating function of [^3] In [4, Section 3], $u(n)$ was denoted by ${\sigma}{\sigma}(n)$ and was called stacks with summits of size $n$. $u(n)$, the number of unimodal sequences of size $n$, is given by (see [4])
@@
Thus $U$ is a mixed false theta function, which a linear combinations of false theta functions multiplied by modular forms. Like Jacobi theta functions fit into the theory of Jacobi form, by introducing an extra Jacobi variable, Nazaroglu and the second author [21, Theorem 1.1] showed that the “modular completion” of a two-variable false theta functions transforms like a Jacobi form.
@@
Various examples of modular and mock modular generating functions arising from partition statistics have been studied extensively in the literature. Classical instances include generating functions of ranks, unimodal sequences, and concave compositions. Here the corresponding generating functions are closely related to modular forms or mock theta functions. In contrast, the restricted overpartition generating functions studied in this paper lead naturally to decompositions involving several different modular objects simultaneously. This highlights a new setting in which mixed modular structures appear in a combinatorial context.
Motivated by the rich modular behaviour exhibited by generating functions of restricted partitions and overpartitions, we investigate several families of overpartitions with parity and overlining restrictions. Such restrictions often lead to generating functions that admit tractable $q$-series representations while still exhibiting interesting modular properties. In this paper, we consider overpartitions of $n$ with the following restrictions:
@@
Here we show that $\overline{P}_{\textup{od}}$ can be written in terms of a false theta function and a mixed mock Maass theta function. [^5] Mixed mock Maass theta functions are linear combinations of products of modular forms and mock Maass theta functions. For this define, [^6] This is a mock Maass theta function by work of Zwegers (see the discussion at the end of Section 6 of [39]). Similar to the case of $\sigma$}, the Fourier coefficients of $W_{1}$ are related to the arithmetic of $\mathbb{Q}(\sqrt{2})$. following Corson, Favero, Leisinger, and Zubairy [25],
@@
A notable feature of the generating functions arising in this work is that they involve simultaneously several types of modular objects. While generating functions in partition theory are often related to a single modular object (e.g. a (mock) modular form), the examples studied here naturally lead to decompositions involving several different modular structures at once. For instance, in Theorem 1.1 we express the generating function Pod(q) as a sum of a mixed mock Maass theta function, a modular form, and a false theta function. Similar phenomena appear for the other generating functions considered in this paper. These examples suggest that natural restrictions on overpartitions can lead to generating functions exhibiting mixed modular behaviour.
@@
Using this, [10, Theorem 6.1], Theorem 1.3, and the facts that $g(0)=0$ and $g(1)=1$, we obtain
@@
We recall another (see [22, p. 347]) generalized [^8] From [22, p. 347], we know that for $k,\ell\in\mathbb{N}$, $\nu(q^{k},q^{\ell};q)$ is a mock theta function in the sense of Ramanujan. third order mock theta function
@@
Remark.
Taken together, our results illustrate that natural restrictions on overpartitions and concave compositions can lead to generating functions exhibiting mixed modular structures. In particular, our examples show that combinations of modular forms, mock theta functions, mock Maass theta functions, and false theta functions arise naturally in combinatorial generating functions.
@@ 2.1. $q$-series identities
Recall the following two identities from [6, equation (2.2.5) and equation (2.2.6)].
@@
Next, we require an identity due to Garvan [29, Theorem 1.2].
@@
Due to Andrews [2, equation (4.4)], we have the following:
@@
In [26, Theorem 1.1], we find the following identity.
@@
From [7, equation (2.3.1)], we have the following lemma.
@@
From [6, Corollary 2.4], we record the following identity.
@@
From [6, Example 10, p. 29], we record the following identity.
@@
We also require identities by Bhoria, Eyyunni, and Maji. The first is (see [16, Theorem 2.1]).
@@
The second identity (see [16, Theorem 2.2]) is the following:
@@
Next, we list the following identity from [14, equation (3.2.6)].
@@
Finally, we require the Rogers–Fine identity [28, equation (14.1)].
@@ 2.2. A Tauberian theorem
We also require a modified version of Ingham’s Tauberian Theorem [31], stated in [20, Theorem 1.1]. For $\Delta\geq 0$ define
@@ 2.3. The Euler–Maclaurin summation formula
We use a variation of the Euler–Maclaurin summation formula, an important tool for computing asymptotic expansions of infinite sums, studied by Zagier [36, Proposition 3] for one-dimensional smooth functions and extended by Jennings-Shaffer, Mahlburg, and the second author [20, Theorem 1.2] for multi-dimensional complex-valued functions. A function $g$, defined over an unbounded domain $D\subset\mathbb{C}$, is of sufficient decay if there exists $\varepsilon>0$ such that $g(w)\ll|w|^{-1-\varepsilon}$ uniformly as $|w|\to\infty$ in $D$. For $0\leq\theta<\frac{\pi}{2}$, we define $D_{\theta}:=\{re^{i\alpha}:r\geq 0\ \text{and}\ |\alpha|\leq\theta\}$.
@@
Next, we consider two-dimensional sums. Throughout we write $\bm{x}=\left(x_{1},x_{2}\right)\in\mathbb{R}^{2}$. A function $f$ defined on an unbounded domain $D\subset\mathbb{C}^{2}$ is of sufficient decay if there exist $\varepsilon_{1},\varepsilon_{2}>0$ such that $f(\bm{x})\ll|w_{1}|^{-1-\varepsilon_{1}}|w_{2}|^{-1-\varepsilon_{2}}$ uniformly as $|w_{1}|+|w_{2}|\to\infty$ in $D$. From [20, Proposition 5.2] (with $r=2$), we have the following.
@@
From [25, Theorem 2.3], we have
@@
Plugging this into (5), using (1.9), using (1.4), [10, equation (1.6)], [28, equation (26.27)], and (1.13), we conclude the proof.∎
@@ 6. Concluding Remarks
The results of this paper suggest several directions for further investigation. In particular, it would be interesting to understand more systematically when restrictions on overpartitions lead to generating functions exhibiting mixed modular structures. Another natural problem is to study whether similar phenomena occur for other partition statistics or colored partition families. The results of this paper suggest several directions for further investigation.
@@
  3. (3)
    The generating functions studied here exhibit combinations of several different modular objects, including modular forms, mock theta functions, mock Maass theta functions, and false theta functions. It would be interesting to understand more systematically when restrictions on overpartitions lead to generating functions with such mixed modular structures.
  4. (4)
    Another direction would be to define and study additional rank or crank-type statistics for the restricted overpartitions considered here and to investigate the modular properties of their generating functions.
  5. (5)
    In this paper, we determine the asymptotic main terms of several counting functions. It would be interesting to obtain more precise asymptotic expansions, possibly using the Circle Method or modular transformations.
  6. (6)
    It would be interesting to see whether the restricted overpartition functions studied here satisfy arithmetic congruences analogous to those known for the classical partition function.